OpenStudy (anonymous):

What is the y-value for f(x) = tan(x) when x = -45°?

Nnesha (nnesha):

replace x by -45

Nnesha (nnesha):

remember: \[\huge\rm \tan = \frac{ \sin }{ \cos }\]

OpenStudy (studygurl14):

Have you tried inputting tan(-45) into a calcualtor?

OpenStudy (jdoe0001):

\(\bf tan(-45^o)=\cfrac{sin(-45^o)}{cos(-45^o)}\qquad \begin{cases} sin(-\theta)=-sin(\theta)\\ cos(-\theta)=cos(\theta) \end{cases} \\ \quad \\ \cfrac{sin(-45^o)}{cos(-45^o)}\implies \cfrac{-sin(45^o)}{cos(45^o)}\implies -\cfrac{sin(45^o)}{cos(45^o)}\implies -tan(45^o)\) check your Unit Circle, or as suggested already, use your calculator